Show that the density of nucleus is 2.3×1017 kgm- 3
1
Expert's answer
2020-06-12T11:12:01-0400

Take sodium for example with mas A=23.

The radius of the nucleus is approximately


"r=r_0A^{1\/3}."

The volume is


"V=\frac{4}{3}\pi r^3."

The mass in kilograms can be obtained from A:

"m=A\cdot1.66\cdot10^{-27}\text{ kg}."

The density is


"\rho=\frac{m}{V}=\frac{3A\cdot1.66\cdot10^{-27}\text{ kg}}{4\pi r_0^3A}=\frac{3\cdot1.66\cdot10^{-27}\text{ kg}}{4\pi r_0^3},\\\space\\
\rho=\frac{3\cdot1.66\cdot10^{-27}\text{ kg}}{4\pi (1.2\cdot10^{-15})^3}=2.29\cdot10^{17}\text{ kg/m}^3."

As we see, the density does not depend on the mass number A.